# FlyPose: Towards Robust Human Pose Estimation From Aerial Views

### v1

@@ Abstract
Unmanned Aerial Vehicles (UAVs) are increasingly deployed in close proximity to humans for applications such as parcel delivery, traffic monitoring, disaster response and infrastructure inspections. Ensuring safe and reliable operation in these human-populated environments demands accurate perception of human poses and actions from an aerial viewpoint. This perspective challenges existing methods with low resolution, steep viewing angles and (self-)occlusion, especially if the application demands realtime feasibile models. We train and deploy FlyPose, a lightweight top-down human pose estimation pipeline for aerial imagery. Through multi-dataset training, we achieve an average improvement of 6.8 mAP in person detection across the test-sets of Manipal-UAV, VisDrone, HIT-UAV as well as our custom dataset. For 2D human pose estimation we report an improvement of 16.3 mAP on the challenging UAV-Human dataset. FlyPose runs with an inference latency of $\scriptstyle\sim$20 milliseconds including preprocessing on a Jetson Orin AGX Developer Kit and is deployed onboard a quadrotor UAV during flight experiments. We also publish FlyPose-104, a small but challenging aerial human pose estimation dataset, that includes manual annotations from difficult aerial perspectives: https://github.com/farooqhassaan/FlyPose.
@@ 1 Introduction
With rising levels of automation in commercial unmanned aerial vehicles (UAVs, also commonly referred to as drones), new applications such as parcel delivery, search-and-rescue, infrastructure inspection and urban traffic monitoring are being introduced. Their operation in human-populated spaces [2] makes the robust perception of human presence and behaviors indispensible. In both human-human and human-machine interaction, the configuration of body joints conveys non-verbal cues for understanding body language, physical movement and intent. The task of human pose estimation (HPE) represents a widely-used foundational module to extract these cues from images to enable skeleton-based downstream tasks such as action recognition, pose-based tracking, motion prediction, engagement detection or interaction via hand gestures. Based on the insights gained from these tasks, the UAV can then respond appropriately based on its own behavior model. The majority of applications for HPE cannot assume a ground-level or shoulder-height camera view where the human subject is seen frontally or laterally, as this viewpoint would enable the observation of distinguishable human body features. Aerial viewpoints tend to occlude pose keypoints of the face or legs, distort body proportions and challenge pose models trained on ground-level images. Therefore, our work explores human pose estimation from an aerial perspective to aid UAV applications in human-populated environments (Figure 1).
@@ 2 Related Work
For Pose Estimation, ViT-based architectures like ViTPose [49] also represent the current state-of-the art on the COCO test-dev set, while examples for lightweight architectures with competitive performance are RTMO [26], RTMPose [18], KAPAO [29] and YOLO-Pose [28]. Nevertheless, their accuracy drops notably when applied to UAV imagery, due to viewpoint shifts and scale variation, even when not accounting for small objects. Top-down views like in CMU Panoptic [19], ITOPS [13, 11] and PoseFES [53] are rarely explored outside controlled conditions. There have been very few recent works [20, 17] that try to address the challenge of aerial human pose estimation from UAVs. In [20], they try to evaluate the performance of out-of-the-box pose estimators on various datasets and gauge their performance on edge devices. The authors of [17] instead attempt to refine an estimated pose to resemble poses from an aerial poses’ latent codebook. A small number of aerial datasets like Aerial Gait [31] or the synthethic dataset in Airpose [39] exist, but they are limited in variety and scale. Airpose attempts multi-view 3D human pose and shape estimation but the used methods still rely on 2D pose priors from OpenPose and AlphaPose which were only finetuned for slightly elevated views from the side without top-views or covering larger distances. UAV-Human [22] is the largest aerial pose dataset so far, offering annotated human keypoints from drone footage in diverse settings.
@@ 3 Methodology
Figure 3: System overview: the bottom illustrates the FlyPose pipeline, where the detector and pose estimation model are trained separately, the top is an example for how the aggregated information can be used for downstream tasks within the drone system for various applications.
@@ 3.2 Evaluation Metrics
For person detection, we use the COCO mean Average Precision (mAP) which averages the Average Precisions across the 0.5-0.95 intersection over union (IoU) thresholds and the Average Recall (AR) for 100 allowed predictions. Since we conduct multi-dataset training, the Average Precision and Recall across all testing sets, weighted by the number of corresponding frames, is reported for reference. We do not report classification scores, since all datasets were filtered to only include one person class. For human pose estimation, we evaluate the model performance using standard COCO mean Average Precision (mAP) for keypoints, similarly averaging across the 0.5-0.95 Object Keypoint Similarity (OKS) thresholds. We use the standard OKS metric by COCO [23] that measures how close predicted keypoints are to the ground-truth, normalized by the person scale (bounding box area) and a per-keypoint tolerance. To measure the feasibility of the models in practice, we measure the latency of the models as the time taken for a single forward pass through the model with a batch size of 1. The latency of the person detection and pose estimation models are reported separately, with the expectation of the latency of pose estimation to increase for multiple persons because of the top-down pose estimation. Although this can be offset by batching person detections together and incurring a cost of preparing the batch, further experimentation is necessary. The inference latency for the models is measured on a Mobile GPU A6000 as well as on the Jetson Orin AGX Developer Kit with 32 GB RAM. Note that on the Jetson, the models were converted to FP32 TensorRT engines for lower latency.
@@ 4.1 Aerial Person Detection
As a first step, RT-DETRv2-S with pretrained weights from COCO [23] and Objects365 [40] is finetuned for 60 epochs on the VisDrone2019-DET dataset [9] after removing all classes other than “pedestrian” and “person”, which are combined into a single person class. After the finetuning, the model performs 7.1 mAP higher on average across all of the used test-sets. This first step is standard practice but VisDrone only features imagery from nearby the city of Tianjin, therefore we proceed to gather additional aerial datasets featuring a person class to further improve the model’s generalization ability outside urban scenarios. We add the original train/val split of eight additional aerial object detection datasets, resulting in 66849 additional images for training and 21164 for validation besides VisDrone2019-DET. To give a brief overview, we describe them here: SeasDronesSea [44] and Heridal [3] represent RGB aerial imagery for search-and-rescue (SAR) purposes, where the former has maritime and the latter hilly, mediterranean terrain as background. VTSAR [56], DroneRGBT [57], VTUAV-det [55] feature both RGB and thermal imagery from UAVs for training, while HIT-UAV [43] is the only pure thermal dataset with a separate testing split. Manipal-UAV [1] features person detection at altitudes from 10 to 50 meters with complex backgrounds and TinyPerson [54] is the most challenging dataset for our approach, since it focusses purely on dense clusters of tiny people from an aerial perspective. Like with VisDrone, only the person-related classes are kept in the annotations to test if the model learns robust features although presented with both aerial frontal and top-views. Rather than inspecting their individual performance, we combine them into one dataset together with VisDrone and the resulting test set performance is reported in the row “Multi-Dataset” in Tab. 1. Note that we did not exclude thermal imagery because for night-time applications like SAR, the model has to be able to predict on both modalities. As a result, the detector gains another 6.78 mAP from the previous iteration on average but mainly caused by the introduction of the thermal imagery and the resulting improvement in the HIT-UAV test-set.
@@
Since the frontal perspectives within the COCO dataset are still useful in frontal low-altitude perspectives and the initial mAP on the COCO validation set decreased after finetuning on aerial data, we reintroduce the COCO-Person train-val split into the finetuning to mitigate this effect. This step slightly decreased the average mAP on the aerial test-sets by 0.14 mAP but improved the average AR by 1.01. Finally, to improve the localization on smaller objects, Normalized Wasserstein Distance Loss [48] was introduced for the final training which had only a neglible impact overall but it did improve both mAP and AR on our FlyPose-104 test set. Although individual training steps led to the highest increase of mAP on one test-set (for example, only finetuning on VisDrone had the best results on the respective test set), our multi-dataset version of RT-DETRv2-S generalizes robustly across all used aerial test sets (examples in Figure 5). To establish the model’s realtime feasibility, we measured 13 milliseconds
@@ 4.2 Aerial Pose Estimation
ViTPose-S, in particular, achieves the fastest performance on the Jetson Orin with an inference time of just 6.54 ms when converted to a TensorRT FP32 engine, making it the most suitable candidate for our onboard system. Given the realtime demands of various aerial applications, ViTPose-S is selected as the core HPE model for FlyPose. A qualitative example of the full FlyPose pipeline showing detection and pose estimation outputs is presented in Figure 6. In combination, our final FlyPose model consisting of RT-DETRv2-S and ViTPose-S has an inference latency of only 19.54 miliseconds on a Jetson Orin AGX developer kit for a single image forward pass, excluding pre- or post-processing.
@@ 4.3 Deployment onboard the UAV
To validate FlyPose in real flight with a UAV, the Jetson Orin AGX developer kit with 32GB of RAM and a multi-spectral gimbal camera were mounted onto a commercial quadrotor UAV with a maximum take-off weight of 35kg, and connected via Ethernet to receive images from the camera’s Full-HD RTSP stream. We then run FlyPose on the embedded system to predict poses in order to estimate where a person is pointing in a simulated cargo pickup task. Figure 7 shows a snapshot from one of the flight experiments, where the left image is a view of the onboard camera and the right shows the UAV itself from an external perspective.
@@ 5 Discussion
In the person detection stage of our pipeline, we learned that while training on aerial imagery significantly improved the model’s predictions from overhead views, lightweight detectors like RT-DETRv2 still face challenges with small objects, even after training on datasets like TinyPerson. Similar to human perception, in order to distinguish small scale objects from background noise, temporal motion features are sometimes essential. Mixing top-down aerial views with frontal imagery from datasets like COCO frequently lead to the model omitting an extended arm when viewing the person from the top or falsely including a background object that resembles an arm. Especially, when the person is wearing darker clothing, the boundary between the persons shadow and the feet for example can be visually ambiguous. Cluttered backgrounds with unfamiliar elements still frequently led to false positives, whereas increasing the confidence threshold also prevented distant objects from being picked up. Although top-down and frontal views, as well as RGB and thermal imagery showcase different visual features, we chose not to separate them into distinct classes. Instead we train on a merged multi-modal dataset, where the person class still represents the same object category, and an applied model should be able to handle viewpoint and color variations. While our detection precision is not on par with larger scale state-of-the-art models on each individual dataset, RT-DETRv2-S still demonstrates good generalization across multiple test sets, views and modalities at a small memory footprint and is realtime feasibile on an embedded system.
@@
Examples of the pose results of our FlyPose model on the UAV-Human, Manipal, Okutama, FlyPose-104 and VisDrone datasets are shown in Figure 8. The COCO-pretrained ViTPose-S baseline model frequently struggles with discerning self-occluded keypoints, these are better estimated by our finetuned model. It can be observed that it performs relatively better on low resolution patches as well. Cases where our model still struggles are with persons strongly camouflaged by the background or when detected patches contain multiple people in close proximity. Qualitative samples of thermal imagery were also inspected, but while the features did generalize to a degree, the poses were not as reliable as on aerial RGB images. Due to the lack of color information at lower resolutions, left-right keypoint identity swaps occurred more frequently when it was visually ambigous which direction the person is facing.
We also observe that the keypoint heatmaps show weak confidence for joint detections in low-resolution patches, as they are frequently cut-off by the used confidence threshold of 0.4 for keypoint estimates. Nevertheless, for individuals at greater distances, person detection may suffice for most applications. Our additional evaluation results on the UAV-Human dataset revealed a consistent drop in average Object Keypoint Similarity (OKS) scores for facial keypoints such as nose, eyes and ears, compared to body keypoints like shoulders, elbows, wrists, hips, knees and feet. This suggests that facial landmarks are particularly challenging to localize from aerial viewpoints, likely due to occlusions and limited visibility. Although more analysis on the impact of high-person-count images on the system needs to be conducted, one way to deal with this could be an informed selection strategy that only performs keypoint estimation on relevant detected people depending on the current task of the UAV. Finally, despite the current challenges in aerial human pose estimation, FlyPose is a first step to more robustly recognize human poses and perform downstream tasks like gesture and action recognition from aerial views, directly onboard the UAV.

### v2

@@ Abstract
Unmanned Aerial Vehicles (UAVs) are increasingly deployed in close proximity to humans for applications such as parcel delivery, traffic monitoring, disaster response and infrastructure inspections. Ensuring safe and reliable operation in these human-populated environments demands accurate perception of human poses and actions from an aerial viewpoint. This perspective challenges existing methods with low resolution, steep viewing angles and (self-) occlusion, especially if the application demands realtime feasible models. We train and deploy FlyPose, a lightweight top-down human pose estimation pipeline for aerial imagery. Through multi-dataset training, we achieve an average improvement of 6.8 mAP in person detection across the test-sets of Manipal-UAV, VisDrone, HIT-UAV as well as our custom dataset. For 2D human pose estimation we report an improvement of 16.3 mAP on the challenging UAV-Human dataset. FlyPose runs with an inference latency of $\scriptstyle\sim$20 milliseconds including preprocessing on a Jetson Orin AGX Developer Kit and is deployed onboard a quadrotor UAV during flight experiments. We also publish FlyPose-104, a small but challenging aerial human pose estimation dataset, that includes manual annotations from difficult aerial perspectives: https://github.com/farooqhassaan/FlyPose.
@@ 1 Introduction
With rising levels of automation in commercial unmanned aerial vehicles (UAVs, also commonly referred to as drones), new applications such as parcel delivery, search-and-rescue, infrastructure inspection and urban traffic monitoring are being introduced. Their operation in human-populated spaces [2] makes the robust perception of human presence and behaviors indispensable. In both human-human and human-machine interaction, the configuration of body joints conveys non-verbal cues for understanding body language, physical movement and intent. The task of human pose estimation (HPE) represents a widely-used foundational module to extract these cues from images to enable skeleton-based downstream tasks such as action recognition, pose-based tracking, motion prediction, engagement detection or interaction via hand gestures. Based on the insights gained from these tasks, the UAV can then respond appropriately based on its own behavior model. The majority of applications for HPE cannot assume a ground-level or shoulder-height camera view where the human subject is seen frontally or laterally, as this viewpoint would enable the observation of distinguishable human body features. Aerial viewpoints tend to occlude pose keypoints of the face or legs, distort body proportions and challenge pose models trained on ground-level images. Therefore, our work explores human pose estimation from an aerial perspective to aid UAV applications in human-populated environments (Figure 1).
@@ 2 Related Work
For Pose Estimation, ViT-based architectures like ViTPose [49] also represent the current state-of-the art on the COCO test-dev set, while examples for lightweight architectures with competitive performance are RTMO [26], RTMPose [18], KAPAO [29] and YOLO-Pose [28]. Nevertheless, their accuracy drops notably when applied to UAV imagery, due to viewpoint shifts and scale variation, even when not accounting for small objects. Top-down views like in CMU Panoptic [19], ITOPS [13, 11] and PoseFES [53] are rarely explored outside controlled conditions. There have been very few recent works [20, 17] that try to address the challenge of aerial human pose estimation from UAVs. In [20], they try to evaluate the performance of out-of-the-box pose estimators on various datasets and gauge their performance on edge devices. The authors of [17] instead attempt to refine an estimated pose to resemble poses from an aerial poses’ latent codebook. A small number of aerial datasets like Aerial Gait [31] or the synthetic dataset in Airpose [39] exist, but they are limited in variety and scale. Airpose attempts multi-view 3D human pose and shape estimation but the used methods still rely on 2D pose priors from OpenPose and AlphaPose which were only finetuned for slightly elevated views from the side without top-views or covering larger distances. UAV-Human [22] is the largest aerial pose dataset so far, offering annotated human keypoints from drone footage in diverse settings.
@@ 3 Methodology
Figure 3: System overview: the bottom illustrates the FlyPose pipeline, where the detector and pose estimation model are trained separately, and the top is an example of how the aggregated information can be used for downstream tasks within the drone system for various applications.
@@ 3.2 Evaluation Metrics
For person detection, we use the COCO mean Average Precision (mAP) which averages the Average Precisions across the 0.5-0.95 intersection over union (IoU) thresholds and the Average Recall (AR) for 100 allowed predictions. Since we conduct multi-dataset training, the Average Precision and Recall across all testing sets, weighted by the number of corresponding frames, are reported for reference. We do not report classification scores, since all datasets were filtered to only include one person class. For human pose estimation, we evaluate the model performance using standard COCO mean Average Precision (mAP) for keypoints, similarly averaging across the 0.5-0.95 Object Keypoint Similarity (OKS) thresholds. We use the standard OKS metric by COCO [23] that measures how close predicted keypoints are to the ground-truth, normalized by the person scale (bounding box area) and a per-keypoint tolerance. To measure the feasibility of the models in practice, we measure the latency of the models as the time taken for a single forward pass through the model with a batch size of 1. The latency of the person detection and pose estimation models are reported separately, with the expectation of the latency of pose estimation to increase for multiple persons because of the top-down pose estimation. Although this can be offset by batching person detections together and incurring a cost of preparing the batch, further experimentation is necessary. The inference latency for the models is measured on a Mobile GPU A6000 as well as on the Jetson Orin AGX Developer Kit with 32 GB RAM. Note that on the Jetson, the models were converted to FP32 TensorRT engines for lower latency.
@@ 4.1 Aerial Person Detection
As a first step, RT-DETRv2-S with pretrained weights from COCO [23] and Objects365 [40] is finetuned for 60 epochs on the VisDrone2019-DET dataset [9] after removing all classes other than “pedestrian” and “person”, which are combined into a single person class. After the finetuning, the model performs 7.1 mAP higher on average across all of the used test-sets. This first step is standard practice but VisDrone only features imagery from nearby the city of Tianjin, therefore we proceed to gather additional aerial datasets featuring a person class to further improve the model’s generalization ability outside urban scenarios. We add the original train/val split of eight additional aerial object detection datasets, resulting in 66849 additional images for training and 21164 for validation besides VisDrone2019-DET. To give a brief overview, we describe them here: SeasDronesSea [44] and Heridal [3] represent RGB aerial imagery for search-and-rescue (SAR) purposes, where the former has maritime and the latter hilly, Mediterranean terrain as background. VTSAR [56], DroneRGBT [57], VTUAV-det [55] feature both RGB and thermal imagery from UAVs for training, while HIT-UAV [43] is the only pure thermal dataset with a separate testing split. Manipal-UAV [1] features person detection at altitudes from 10 to 50 meters with complex backgrounds and TinyPerson [54] is the most challenging dataset for our approach, since it focuses purely on dense clusters of tiny people from an aerial perspective. Like with VisDrone, only the person-related classes are kept in the annotations to test if the model learns robust features although presented with both aerial frontal and top-views. Rather than inspecting their individual performance, we combine them into one dataset together with VisDrone and the resulting test set performance is reported in the row “Multi-Dataset” in Tab. 1. Note that we did not exclude thermal imagery because for nighttime applications like search-and-rescue, the model has to be able to predict on both modalities. As a result, the detector gains another 6.78 mAP from the previous iteration on average but mainly caused by the introduction of the thermal imagery and the resulting improvement in the HIT-UAV test-set.
@@
Since the frontal perspectives within the COCO dataset are still useful in frontal low-altitude perspectives and the initial mAP on the COCO validation set decreased after finetuning on aerial data, we reintroduce the COCO-Person train-val split into the finetuning to mitigate this effect. This step slightly decreased the average mAP on the aerial test-sets by 0.14 mAP but improved the average AR by 1.01. Finally, to improve the localization on smaller objects, Normalized Wasserstein Distance Loss [48] was introduced for the final training which had only a negligible impact overall but it did improve both mAP and AR on our FlyPose-104 test set. Although individual training steps led to the highest increase of mAP on one test-set (for example, only finetuning on VisDrone had the best results on the respective test set), our multi-dataset version of RT-DETRv2-S generalizes robustly across all used aerial test sets (examples in Figure 5). To establish the model’s realtime feasibility, we measured 13 milliseconds
@@ 4.2 Aerial Pose Estimation
ViTPose-S, in particular, achieves the fastest performance on the Jetson Orin with an inference time of just 6.54 ms when converted to a TensorRT FP32 engine, making it the most suitable candidate for our onboard system. Given the realtime demands of various aerial applications, ViTPose-S is selected as the core HPE model for FlyPose. A qualitative example of the full FlyPose pipeline showing detection and pose estimation outputs is presented in Figure 6. In combination, our final FlyPose model consisting of RT-DETRv2-S and ViTPose-S has an inference latency of only 19.54 milliseconds on a Jetson Orin AGX developer kit for a single image forward pass, excluding pre- or post-processing.
@@ 4.3 Deployment onboard the UAV
To validate FlyPose in real flight with a UAV, the Jetson Orin AGX developer kit with 32GB of RAM and a multi-spectral gimbal camera were mounted onto a commercial quadrotor UAV with a maximum take-off weight of 35kg, and connected via Ethernet to receive images from the camera’s Full-HD RTSP stream. We then ran FlyPose on the embedded system to predict poses in order to estimate where a person is pointing in a simulated cargo pickup task. Figure 7 shows a snapshot from one of the flight experiments, where the left image is a view of the onboard camera and the right shows the UAV itself from an external perspective.
@@ 5 Discussion
In the person detection stage of our pipeline, we learned that while training on aerial imagery significantly improved the model’s predictions from overhead views, lightweight detectors like RT-DETRv2 still face challenges with small objects, even after training on datasets like TinyPerson. Similar to human perception, in order to distinguish small scale objects from background noise, temporal motion features are sometimes essential. Mixing top-down aerial views with frontal imagery from datasets like COCO frequently led to the model omitting an extended arm when viewing the person from the top or falsely including a background object that resembles an arm. Especially, when the person is wearing darker clothing, the boundary between the persons shadow and the feet for example can be visually ambiguous. Cluttered backgrounds with unfamiliar elements still frequently led to false positives, whereas increasing the confidence threshold also prevented distant objects from being picked up. Although top-down and frontal views, as well as RGB and thermal imagery showcase different visual features, we chose not to separate them into distinct classes. Instead we train on a merged multi-modal dataset, where the person class still represents the same object category, and an applied model should be able to handle viewpoint and color variations. While our detection precision is not on par with larger scale state-of-the-art models on each individual dataset, RT-DETRv2-S still demonstrates good generalization across multiple test sets, views and modalities at a small memory footprint and is realtime feasibile on an embedded system.
@@
Examples of the pose results of our FlyPose model on the UAV-Human, Manipal, Okutama, FlyPose-104 and VisDrone datasets are shown in Figure 8. The COCO-pretrained ViTPose-S baseline model frequently struggles with discerning self-occluded keypoints, these are better estimated by our finetuned model. It can be observed that it performs relatively better on low resolution patches as well. Cases where our model still struggles are with persons strongly camouflaged by the background or when detected patches contain multiple people in close proximity. Qualitative samples of thermal imagery were also inspected, but while the features did generalize to a degree, the poses were not as reliable as on aerial RGB images. Due to the lack of color information at lower resolutions, left-right keypoint identity swaps occurred more frequently when it was visually ambiguous which direction the person is facing.
We also observe that the keypoint heatmaps show weak confidence for joint detections in low-resolution patches, as they are frequently cut-off by the used confidence threshold of 0.4 for keypoint estimates. Nevertheless, for individuals at greater distances, person detection may suffice for most applications. Our additional evaluation results on the UAV-Human dataset revealed a consistent drop in average OKS scores for facial keypoints such as nose, eyes and ears, compared to body keypoints like shoulders, elbows, wrists, hips, knees and feet. This suggests that facial landmarks are particularly challenging to localize from aerial viewpoints, likely due to occlusions and limited visibility. Although more analysis on the impact of high-person-count images on the system needs to be conducted, one way to deal with this could be an informed selection strategy that only performs keypoint estimation on relevant detected people depending on the current task of the UAV. Finally, despite the current challenges in aerial human pose estimation, FlyPose is a first step to more robustly recognize human poses and perform downstream tasks like gesture and action recognition from aerial views, directly onboard the UAV.